You are a manager of Nvidia and your only significant competitor in the mainstream graphics card market is the ATI subsidiary of Advanced Micro Devices. ATI and Nvidia both expect to select their output of the next generation of graphics card in July of next year. Nvidia’s graphics cards and ATI’s graphics cards are indistinguishable to consumers. The market demand for graphics cards is Q = 3 – 0.25 P (with price P in dollars) and both firms used to produce at a constant unit cost of $4. However, the engineering department at Nvidia estimates that it can reduce their own marginal cost to $3 if they develop a new procedure. That new procedure is estimated to cost $2 to develop. Should Nvidia develop that new procedure?
Answer #1

Demand : Q = 3 - 0:25 p ::P : 12-16 :12 - 192-492 unit cost : MC : AC = 7 Profits : 7,89, -79,92-191 - T2 = 892-492 49,92 FOCThe profit decreases with innovation, hence, the firm must not develop the new procedure.